# Systematic biases from the exclusion of higher harmonics in parameter estimation on LISA binaries

### v1

@@ II.3 Predicting biases with direct likelihood optimization
In Sec. V we will confirm the finding of Ref. [40]: the biases due to neglecting higher-order modes are not always well-approximated via the CV formalism, in particular for more massive events. Given this result, it is desirable to explore other methods of predicting the biases we see in PE. In this section, we leverage the fact that the systematic bias we consider here is, by definition, the difference between the “true” (or injected) parameters and the parameters evaluated at the maximum likelihood given our best-fit waveform. Studying this bias, therefore, in principle reduces to a straightforward optimization problem. Indeed, in theory, performing a rigorous sampling from the posteriors might be considered excessive if one is only interested in studying the distance between the maximum likelihood point and the injection (assuming that the maximum likelihood point can be found without a sampler).
We will see in Sec. V that, just as expected, one can indeed infer the systematic bias with remarkable success when simply optimizing the likelihood directly. In this work, we use the Nelder-Mead algorithm [50] to maximize the likelihood. Essentially, in the Nelder-Mead algorithm, a “simplex” with $n+1$ vertices ($n$ being the number of variables) is introduced in the parameter space. The function value is evaluated at each vertex, and then the simplex systematically rearranges (growing, reflecting, contracting, etc.) such that it moves toward and eventually closes in on the location of the minimum. One appealing characteristic of this algorithm is that it is able to minimize a function of $n$ variables without differentiation.
In this work, we use the implementation of the Nelder-Mead algorithm provided in scipy [51]. We initialize our simplex in $(n+1)$-dimensional space by setting each vertex of the initial simplex at the likelihood evaluated with one parameter set to a value $m\sigma$ away from its injected value, where $\sigma$ is calculated from the diagonal elements of the Fisher covariance matrix. We then minimize the function $-\mathrm{ln}(\mathcal{L})$ (i.e., we find the maximum likelihood). To accommodate the large range in parameter values (considering, for example, that we have masses $\mathcal{O}(10^{5}-10^{6})\,M_{\odot}$ and angles in the range $[-\pi,\pi]$), we rescale parameters by the $\sigma$’s calculated from the Fisher analysis to perform this minimization. During the maximization procedure, we enforce the same prior boundaries on the parameters used in PE.
@@ III Results: Parameter Estimation with lisabeta
We now present PE results across our grid of parameter space, examining how biases arise due to an insufficient number of modes. In this section, we focus on examining the biases in intrinsic parameter recovery. Corresponding discussions for extrinsic parameters are given in Sec. IV. For simplicity, we restrict our analysis to injected signals with non-spinning progenitors ($\chi_{1}=\chi_{2}=0$). We place our events at a distance corresponding to $z=1$. For the rest of the extrinsic parameters (source localization, etc.), which we expect to have a minimal impact on our analysis, we simply take the arbitrary values listed in Table 1.
@@ III.1.2 Events with $q=4$
This result may reflect the fact that, despite higher-order modes being more significant for more asymmetric binaries (larger $q$), with a larger $q$ there is also comparatively more SNR in the inspiral compared to the merger/post-merger, where higher-order modes are most important. This could be a significant balancing effect for this set of lowest-mass events, for which higher modes are pushed to the high-frequency region of lower sensitivity of LISA. Moreover, note that the signal amplitude for a fixed total mass, and thus the SNR, decreases with increasing $q$.
@@ III.1.3 Events with $q=8$
To summarize Secs. III.1.1III.1.3, we see that the biases on the intrinsic parameters for this set of lowest-mass binaries are fairly negligible, and are certainly much less severe than the biases observed for the considerably heavier MBHB system examined in Ref. [40].
@@ III.2 MBHBs with $M=10^{6}\,M_{\odot}$
We now consider events with $M=10^{6}\,M_{\odot}$. We will see that, as expected, the biases from insufficient mode content in the templates are more significant for these more massive events, which have total SNR in the range of $873.1-4105.2$, compared to the much lower SNRs observed in Sec. III.1.
@@ IV Biases introduced in extrinsic parameter recovery
We now observe how changing the mode content of the templates affects the recovered posteriors on the extrinsic parameters. The sky localization of events detected by LISA will be crucial for performing follow-up electromagnetic observations of MBHB systems. Previous studies have shown that LISA’s inference of a system’s sky position can be multimodal, and is strongly informed by higher harmonics, particularly at high masses [41]. As a consequence, mismodelling the higher-modes content of the signal can be expected to lead to biases. This contrasts with LVK observations, where most of the information about the sky position comes from triangulating the time of arrival in different detectors, which is mostly independent of the physical content of the waveforms. We attempt to understand how biases can change with differing waveform mode content.
@@
    The extrinsic parameters for all the $10^{5}\,M_{\odot}$ events with either $\iota=\pi/3$ or $\iota\approx\pi/2$ are recovered with very little bias for all mode configurations, and only degenerate in $\varphi$ and $\psi_{L}$ for the quadrupole-only templates, as expected. An example is given in the top left panel of Fig. 10 ($210^{5}\,M_{\odot},q=4,\iota\approx\pi/$ 2). The recovery with only the (2, 2) mode is noticeably worse for the nearly face-on events with this total mass (compare top left and right panels of Fig. 10). We have observed that, for a nearly face-on, equal-mass system with this lower total mass, higher-order modes become barely distinguishable, so that adding them to the waveform template does not change the PE significantly (but we do not show the corresponding plot for brevity).
@@
    While the degeneracy along constant lines of $\varphi\,+\,\psi_{L}$ and $\varphi\,-\,\psi_{L}$ noted in Ref. [41] is evident in nearly all posteriors obtained with (2, 2)-mode only templates, this degeneracy is consistently lifted with the inclusion of even just one subdominant mode. This can be seen in any of the four panels of Fig. 10.
@@
    The constraints on the source localization, i.e., the LISA-frame ecliptic longitude and latitude ($\lambda_{L},\beta_{L}$) (among other parameters), improve drastically with the inclusion of one additional mode beyond the (2, 2). This behavior is also noted in Ref. [41].
@@
    Similarly, the distance-inclination degeneracy commonly observed with quadrupole-only templates is broken by the presence of even one additional mode. This is expected, given the different dependence of the different angular harmonics on the inclination angle.
@@ V.2 Using direct likelihood optimization to set boundaries on unbiased parameter space
Having determined that directly optimizing the likelihood is the better method of cheaply estimating systematic biases due to neglecting higher-order modes, we proceeded to make the heatmap shown in the introduction (Fig. 1) using this method. For each point of fixed total mass and mass ratio, we compute the systematic biases, $\Delta_{\rm{sys}}\theta^{i}$, using the direct likelihood optimization method. To account for the fact that the optimization algorithm occasionally fails, we draw a random redshift value from $z_{\rm{opt}}\in[1,10]$ until a redshift is found at which 15 steps of the optimization algorithm can be successfully completed. (Note: 15 steps were generally completed with the very first random redshift that was drawn, consistent with the good performance of optimization via the Nelder-Mead algorithm discussed in Sec. II.3). We stress that there is nothing physically significant about randomly drawing redshift values here; it is merely a convenient way to initialize the optimization process multiple times when necessary, that is, in case it occasionally fails. We find that there are some regions (in particular, around $4\leq q\leq 7$) in which the optimization algorithm seems more sensitive to initial conditions. To mitigate this effect, we continue looking for values of $z_{\rm{opt}}$ at which the optimization is successfully completed until we find multiple initializations (values of $z_{\rm{opt}}$) which return the same values for the systematic biases. We then take such a systematic bias estimation to be “good”.
@@
Figure 15: Same as Fig. 14 but for nearly-edge-on systems. We plot the minimum redshift for unbiased parameter recovery without the (3, 2) mode. Plots are shown for parameters [$\mathcal{M}_{c},q,\chi_{+},\chi_{-}$] on the left, and for the sky localization and distance on the right.
@@ VI Extrapolating the mode content: approximating $\ell\geq 4$ waveforms
In the above results, we see clearly that parameter recovery for a GW signal with radiation modeled in five modes can be significantly biased even when neglecting only one higher-order mode. Given this result, we cannot assume that the biases cease here; that is, we are not guaranteed that biases will be negligible when injecting a signal with six modes and recovering with five or less, injecting with seven modes and recovering with six or less, etc., particularly when we look at potentially even higher-mass events, (i.e., the “golden” MBHB LISA sources with total mass $\sim 10^{7}\,M_{\odot}$ or higher [17]), which will have both a larger overall SNR and more merger/ringdown-dominated signals, where higher-order modes become more important. While a thorough investigation of such effects is only possible with waveform templates that accurately model gravitational radiation in higher angular harmonics for a range of BBH configurations, here we show that with rough approximations to such waveforms motivated by PN theory, we can already begin to see the effects of neglecting the modes that are currently absent in waveform approximants.
@@ VI.1.1 Approximate amplitudes
To construct rough approximations to higher-order mode amplitudes, we follow the procedure used in Ref. [37] to construct higher-order mode templates for searches of the LVK O3 data. This is essentially the quadrupole mapping procedure used to construct the IMRPhenomHM waveform family [52]. In this procedure, higher-order mode amplitudes are constructed by rescaling the (2, 2) amplitude by the lowest-order PN amplitude terms for each higher multipole. These terms are taken from Sec. 3 of Ref. [53] (with updated versions for $\ell\leq 4$ terms in App. E of Ref. [42]). Explicitly, the expressions for the first five ($\ell=m$) terms are
@@
where $\eta$ is the symmetric mass ratio, $\eta=m_{1}m_{2}/M^{2}$ (with $M=m_{1}+m_{2}$), and the $V_{k}$ are frequency-domain PN parameters given by $V_{k}(f)=[2\pi Mf/k]^{1/3}$. These $H_{\ell m}$’s in PN theory are related to the gravitational waveform by the following equations (see Eqs. (9)–(11) of Ref. [53]):
$h(\theta,\phi)$ $=$ $\sum^{+\infty}_{\ell=2}\sum^{\ell}_{m=-\ell}h_{\ell m}\,{}_{-2}Y_{\ell m}(\theta,\phi)$ (7)
@@
Given this information, the authors of Ref. [37] then construct effective amplitude equations for higher-order modes, given the full waveform for the quadrupole, by dividing both the angular dependence and the $H_{\ell m}$’s of each higher-order mode by that of the quadrupole, i.e.,
@@
Note that compared to Eq. (11) of Ref. [37], we have the additional expressions for $\ell=m=5$ and $\ell=m=6$.
In this work, we are interested in reconstructing modes with $\ell\geq 5$, at variance with Ref. [37], which focused on building only the (3, 3) and (4, 4) modes from the quadrupole. In this case, it is interesting to ask whether more accurate higher-order mode waveforms can be built by approximating $\ell\geq 5$ modes from the (3, 3) or (4, 4) modes, rather than the quadrupole. To investigate this, we additionally compute the following terms:
@@
In the top panel of Fig. 16, we plot the approximate amplitudes we obtained for $h_{55}$ and $h_{66}$, rescaling from both the (2, 2) (dashed lines) and (4, 4) (dot-dashed lines) waveforms. The system plotted in Fig. 16 has total mass and mass ratio $M=10^{6}\,M_{\odot}$ and $q=8$. For comparison, we show in the bottom panel how the approximate (3, 3) and (4, 4) amplitude expressions rescaled from the quadrupole (dashed lines) compare to the “exact” IMRPhenomXHM amplitudes (solid lines). This bottom panel is essentially a reproduction of Fig. 6 of Ref. [37], but for a much heavier BBH system. Compared to the top panel of Fig. 6 of Ref. [37], we also show an approximate (4, 4) waveform that is rescaled from the (3, 3) IMRPhenomXHM waveform (dotted yellow line).
Figure 16: Neglecting higher-order modes ($\ell\geq 5$) which are not currently modeled within the IMRPhenomXHM approximant could introduce biases in PE. We therefore approximately model these modes as shown here, and then test them in a PE run in Fig. 18. Top: We model the (5, 5) and (6, 6) modes by rescaling the IMRPhenomXHM modes for (2, 2) and (4, 4) using the PN-motivated relations in Eqs. (VI.1.1) and (VI.1.1). Bottom: As a test of our rescaling method, we show that modeling (3, 3) and (4, 4) using just the (2, 2) IMRPhenomXHM waveform gives fairly accurate results. All mode amplitudes are plotted for an event with $M=10^{6}\,M_{\odot},q=8$. Dashed, dotted and dot-dashed lines correspond to approximate amplitudes obtained by rescaling the “exact” amplitudes for the (2, 2), (3, 3), and (4, 4) modes, respectively.

### v2

@@ II.3 Predicting biases with direct likelihood optimization
In Sec. V we will confirm the finding of Ref. [Pitte:2023ltw]: the biases due to neglecting higher-order modes are not always well-approximated via the CV formalism, in particular for more massive events. Given this result, it is desirable to explore other methods of predicting the biases we see in PE. In this section, we leverage the fact that the systematic bias we consider here is, by definition, the difference between the “true” (or injected) parameters and the parameters evaluated at the maximum likelihood given our best-fit waveform. Studying this bias, therefore, in principle reduces to a straightforward optimization problem. Indeed, in theory, performing a rigorous sampling from the posteriors might be considered excessive if one is only interested in studying the distance between the maximum likelihood point and the injection (assuming that the maximum likelihood point can be found without a sampler).
We will see in Sec. V that, just as expected, one can indeed infer the systematic bias with remarkable success when simply optimizing the likelihood directly. In this work, we use the Nelder-Mead algorithm [Nelder:1965zz] to maximize the likelihood. Essentially, in the Nelder-Mead algorithm, a “simplex” with $n+1$ vertices ($n$ being the number of variables) is introduced in the parameter space. The function value is evaluated at each vertex, and then the simplex systematically rearranges (growing, reflecting, contracting, etc.) such that it moves toward and eventually closes in on the location of the minimum. One appealing characteristic of this algorithm is that it is able to minimize a function of $n$ variables without differentiation.
In this work, we use the implementation of the Nelder-Mead algorithm provided in scipy [2020NatMe..17..261V]. We initialize our simplex in $(n+1)$-dimensional space by setting each vertex of the initial simplex at the likelihood evaluated with one parameter set to a value $m\sigma$ away from its injected value, where $\sigma$ is calculated from the diagonal elements of the Fisher covariance matrix. We then minimize the function $\mathrm{ln}(\mathcal{L})$ (i.e., we find the maximum likelihood). To accommodate the large range in parameter values (considering, for example, that we have masses $\mathcal{O}(10^{5}-10^{6})\,M_{\odot}$ and angles in the range $\pi,\pi]$), we rescale parameters by the $\sigma$’s calculated from the Fisher analysis to perform this minimization. During the maximization procedure, we enforce the same prior boundaries on the parameters used in PE.
@@ III Results: Parameter Estimation with lisabeta
We now present PE results across our grid of parameter space, examining how biases arise due to an insufficient number of modes. In this section, we focus on examining the biases in intrinsic parameter recovery. Corresponding discussions for extrinsic parameters are given in Sec. IV. We remind the reader that while we divide our discussion of intrinsic and extrinsic parameters for a clearer presentation of results, all 11 parameters (intrinsic and extrinsic) were simultaneously sampled in each of our PE runs. For the present, we restrict our analysis to injected signals with non-spinning progenitors ($\chi_{1}=\chi_{2}=0$). We place our events at a distance corresponding to $z=1$. For the rest of the extrinsic parameters (source localization, etc.), which we expect to have a minimal impact on our analysis, we simply take the arbitrary values listed in Table 1.
@@ III.1.2 Events with $q=4$
This result may reflect the fact that, despite higher-order modes being more significant for more asymmetric binaries (larger $q$), with a larger $q$ there is also comparatively more SNR in the inspiral compared to the merger/postmerger, where higher-order modes are most important. This could be a significant balancing effect for this set of lowest-mass events, for which higher modes are pushed to the high-frequency region of lower sensitivity of LISA. Moreover, note that the signal amplitude for a fixed total mass, and thus the SNR, decreases with increasing $q$.
@@ III.1.3 Events with $q=8$
To summarize Secs. III.1.1III.1.3, we see that the biases on the intrinsic parameters for this set of lowest-mass binaries are fairly negligible, and are certainly much less severe than the biases observed for the considerably heavier MBHB system examined in Ref. [Pitte:2023ltw].
@@ III.2 MBHBs with $M=10^{6}\,M_{\odot}$
We now consider events with $M=10^{6}\,M_{\odot}$. We will see that, as expected, the biases from insufficient mode content in the templates are more significant for these more massive events, which have total SNR in the range of 873.14105.2873.1-4105.2, compared to the much lower SNRs observed in Sec. III.1.
@@ IV Biases introduced in extrinsic parameter recovery
We now observe how changing the mode content of the templates affects the recovered posteriors on the extrinsic parameters. The sky localization of events detected by LISA will be crucial for performing follow-up electromagnetic observations of MBHB systems. Previous studies have shown that LISA’s inference of a system’s sky position can be multimodal, and is strongly informed by higher harmonics, particularly at high masses [Marsat:2020rtl]. As a consequence, mismodeling the higher-modes content of the signal can be expected to lead to biases. This contrasts with LVK observations, where most of the information about the sky position comes from triangulating the time of arrival in different detectors, which is mostly independent of the physical content of the waveforms. We attempt to understand how biases can change with differing waveform mode content.
@@
    The extrinsic parameters for all the $10^{5}\,M_{\odot}$ events with either $\iota=\pi/3$ or $\iota\approx\pi/2$ are recovered with very little bias for all mode configurations, and only degenerate in $\varphi$ and $\psi_{L}$ for the quadrupole-only templates, as expected. An example is given in the top left panel of Fig. 10 ($10^{5}\,M_{\odot},q=4,\iota\approx\pi/2$). The recovery with only the (2, 2) mode is noticeably worse for the nearly face-on events with this total mass (compare top left and right panels of Fig. 10). We have observed that, for a nearly face-on, equal-mass system with this lower total mass, higher-order modes become barely distinguishable, so that adding them to the waveform template does not change the PE significantly (but we do not show the corresponding plot for brevity).
@@
    While the degeneracy along constant lines of $\varphi\,+\,\psi_{L}$ and $\varphi\,-\,\psi_{L}$ noted in Ref. [Marsat:2020rtl] is evident in nearly all posteriors obtained with (2, 2)-mode only templates, this degeneracy is consistently lifted with the inclusion of even just one subdominant mode. This can be seen in any of the four panels of Fig. 10.
@@
    The constraints on the source localization, i.e., the LISA-frame ecliptic longitude and latitude ($\lambda_{L},\beta_{L}$) (among other parameters), improve drastically with the inclusion of one additional mode beyond the (2, 2). This behavior is also noted in Ref. [Marsat:2020rtl]. Note that not only is the bias significantly reduced when one mode is added, but the size of the statistical error (i.e., width of the contours) also decreases considerably. The precision of sky localization recovery does not change nearly as significantly between runs with multiple higher-order modes. This could be important information for rapid inference of an MBHB’s sky position for, e.g., attempted electromagnetic follow-ups. We discuss this point further in Appendix D.
@@
    Similarly, the distance-inclination degeneracy commonly observed with quadrupole-only templates is broken by the presence of even one additional mode. This is expected, given the different dependence of the different angular harmonics on the inclination angle. As with the sky position angles, the statistical error on $D_{L}$ also shrinks considerably with the addition of at least one mode beyond the (2, 2).
@@ V.2 Using direct likelihood optimization to set boundaries on unbiased parameter space
Having determined that directly optimizing the likelihood is the better method of cheaply estimating systematic biases due to neglecting higher-order modes, we proceeded to make the heatmap shown in the introduction (Fig. 1) using this method. For each point of fixed total mass and mass ratio, we compute the systematic biases, $\Delta_{\rm{sys}}\theta^{i}$, using the direct likelihood optimization method. To account for the fact that the optimization algorithm occasionally fails, we draw a random redshift value from $z_{\rm{opt}}\in[1,10]$ until a redshift is found at which 15 steps of the optimization algorithm can be successfully completed. (Note: 15 steps were generally completed with the very first random redshift that was drawn, consistent with the good performance of optimization via the Nelder-Mead algorithm discussed in Sec. II.3). We stress that there is nothing physically significant about randomly drawing redshift values here; it is merely a convenient way to initialize the optimization process multiple times when necessary, that is, in case it occasionally fails. We find that there are some regions (in particular, around $4\leq q\leq 7$) in which the optimization algorithm seems more sensitive to initial conditions. To mitigate this effect, we continue looking for values of $z_{\rm{opt}}$ at which the optimization is successfully completed until we find multiple initializations (values of $z_{\rm{opt}}$) which return the same values for the systematic biases. We then take such a systematic bias estimation to be “good.”
@@
Figure 15: Same as Fig. 14 but for nearly-edge-on systems. We plot the minimum redshift for unbiased parameter recovery without the (3, 2) mode. Plots are shown for parameters $\mathcal{M}_{c},q,\chi_{+},\chi_{-}$] on the left, and for the sky localization and distance on the right.
The statistical errors in Figs. 114, and 15 also vary over the parameter space, in addition to the systematic errors. As a consequence, these plots effectively convey the point at which biases become the limiting factor in accurate parameter estimation, rather than showing the absolute size of the biases. Note also that the choice of a $2\sigma$ threshold here is somewhat arbitrary; if we instead use a threshold of $1\sigma$, the upper bound on the minimum redshift range increases to about 10.5 and 19.5 for the inclination angles $\iota=\pi/3$ and $\iota\approx\pi/2$, respectively, while the shapes of the plots look the same.
@@ VI Extrapolating the mode content: approximating $\ell\geq 4$ waveforms
In the above results, we see clearly that parameter recovery for a GW signal with radiation modeled in five modes can be significantly biased even when neglecting only one higher-order mode. Given this result, we cannot assume that the biases cease here; that is, we are not guaranteed that biases will be negligible when injecting a signal with six modes and recovering with five or less, injecting with seven modes and recovering with six or less, etc., particularly when we look at potentially even higher-mass events (i.e., the “golden” MBHB LISA sources with total mass $\sim 10^{7}\,M_{\odot}$ or higher [LISA:2024hlh]), which will have both a larger overall SNR and more merger/ringdown-dominated signals, where higher-order modes become more important. While a thorough investigation of such effects is only possible with waveform templates that accurately model gravitational radiation in higher angular harmonics for a range of BBH configurations, here we show that with rough approximations to such waveforms motivated by PN theory, we can already begin to see the effects of neglecting the modes that are currently absent in waveform approximants.
@@ VI.1.1 Approximate amplitudes
To construct rough approximations to higher-order mode amplitudes, we follow the procedure used in Ref. [Wadekar:2023kym] to construct higher-order mode templates for searches of the LVK O3 data. This is essentially the quadrupole mapping procedure used to construct the IMRPhenomHM waveform family [London:2017bcn]. In this procedure, higher-order mode amplitudes are constructed by rescaling the (2, 2) amplitude by the lowest-order PN amplitude terms for each higher multipole. These terms are taken from Sec. 3 of Ref. [Mishra:2016whh] (with updated versions for $\ell\leq 4$ terms in Appendix E of Ref. [Garcia-Quiros:2020qpx]). Explicitly, the expressions for the first five ($\ell=m$) terms are
@@
where $\eta$ is the symmetric mass ratio, $\eta=m_{1}m_{2}/M^{2}$ (with $M=m_{1}+m_{2}$), and the $V_{k}$ are frequency-domain PN parameters given by $V_{k}(f)=[2\pi Mf/k]^{1/3}$. These $H_{\ell m}$’s in PN theory are related to the gravitational waveform by the following equations (see Eqs. (9)–(11) of Ref. [Mishra:2016whh]):
$h(\theta,\phi)$ $=$ $\sum^{+\infty}_{\ell=2}\sum^{\ell}_{m=-\ell}h_{\ell m}\,{}_{-2}Y_{\ell m}(\theta,\phi)\,$, (7)
@@
Given this information, the authors of Ref. [Wadekar:2023kym] then construct effective amplitude equations for higher-order modes, given the full waveform for the quadrupole, by dividing both the angular dependence and the $H_{\ell m}$’s of each higher-order mode by that of the quadrupole, i.e.,
@@
Note that compared to Eq. (11) of Ref. [Wadekar:2023kym], we have the additional expressions for $\ell=m=5$ and $\ell=m=6$.
In this work, we are interested in reconstructing modes with $\ell\geq 5$, at variance with Ref. [Wadekar:2023kym], which focused on building only the (3, 3) and (4, 4) modes from the quadrupole. In this case, it is interesting to ask whether more accurate higher-order mode waveforms can be built by approximating $\ell\geq 5$ modes from the (3, 3) or (4, 4) modes, rather than the quadrupole. To investigate this, we additionally compute the following terms:
@@
In the top panel of Fig. 16, we plot the approximate amplitudes we obtained for $h_{55}$ and $h_{66}$, rescaling from both the (2, 2) (dashed lines) and (4, 4) (dot-dashed lines) waveforms. The system plotted in Fig. 16 has total mass and mass ratio $M=10^{6}\,M_{\odot}$ and $q=8$. For comparison, we show in the bottom panel how the approximate (3, 3) and (4, 4) amplitude expressions rescaled from the quadrupole (dashed lines) compare to the “exact” IMRPhenomXHM amplitudes (solid lines). This bottom panel is essentially a reproduction of Fig. 6 of Ref. [Wadekar:2023kym], but for a much heavier BBH system. Compared to the top panel of Fig. 6 of Ref. [Wadekar:2023kym], we also show an approximate (4, 4) waveform that is rescaled from the (3, 3) IMRPhenomXHM waveform (dotted yellow line).
Figure 16: Neglecting higher-order modes ($\ell\geq 5$) which are not currently modeled within the IMRPhenomXHM approximant could introduce biases in PE. We therefore approximately model these modes as shown here, and then test them in a PE run in Fig. 18. Top: we model the (5, 5) and (6, 6) modes by rescaling the IMRPhenomXHM modes for (2, 2) and (4, 4) using the PN-motivated relations in Eqs. (VI.1.1) and (10). Bottom: as a test of our rescaling method, we show that modeling (3, 3) and (4, 4) using just the (2, 2) IMRPhenomXHM waveform gives fairly accurate results. All mode amplitudes are plotted for an event with $M=10^{6}\,M_{\odot},q=8$. Dashed, dotted and dot-dashed lines correspond to approximate amplitudes obtained by rescaling the “exact” amplitudes for the (2, 2), (3, 3), and (4, 4) modes, respectively.